OpenStudy (anonymous):

Which of the following is equivalent to log8 (24)? I put log(8)/log(24), but it was marked wrong. I used the change of base formula to do it. I don't understand this question. Could anyone explain how to do this?

OpenStudy (anonymous):

It looks like this\[\log _{8} 24\] And then my answer was \[\frac{ \log _{8} }{ \log _{24} }\]

OpenStudy (anonymous):

\[FORMULA: \log_{a}b=\frac{ \log_{c} b }{ \log_{c} a } \] Assuming that the changed base is c=8=> \[\log_{8} 24=\frac{ \log_{8} 24 }{ \log_{8} 8}\] which is true considering the fact that \[\log_{8} 8=1\]

OpenStudy (anonymous):

So.. it's \[\frac{ \log 24 }{ \log 8 }\]? Also, a correction on my above statement. My answer WAS \[\frac{ \log 8 }{ \log 24 }\] And there were no logs with the base 8 in any of the answers. \[\log _{8}\] Did I just flip the formula by accident?

OpenStudy (anonymous):

when you write \[\log_{} 24 \] you have a log without base (a nonsens)

OpenStudy (anonymous):

Here, this is what the actual question looks like

OpenStudy (anonymous):

Wouldn't the log have a base of ten if nothings's written as the base?

OpenStudy (anonymous):

no it's not nothing is actually noted lg

OpenStudy (anonymous):

but in this circumstances the correct answer is the first one

OpenStudy (anonymous):

I can give u the explanation

OpenStudy (anonymous):

So basically all of the possible answers are wrong? Am I just learning a really watered down form of logs, maybe? I'm using FLVS, so I wouldn't be surprised

OpenStudy (anonymous):

Sure. This base stuff is seriously confusing me.

OpenStudy (anonymous):

now pay attention to my explication

OpenStudy (anonymous):

you have \[\log_{8} 24\] which can be written with the base change formula as; \[\log_{8} 24=\frac{ \log_{10}24 }{ \log_{10} 8}\]

OpenStudy (anonymous):

they just omitted the 10 from the base using another notation but in principle is right

OpenStudy (anonymous):

Okay. Now, how come that automatic ten appears when you do your change of base formula?

OpenStudy (anonymous):

\[FORMULA:\log_{a}b=\frac{ \log_{c} b }{ \log_{c} a} \] in our case we have: a=8(the original base); b=24(the exponent of the log); c=10(the changed base);

OpenStudy (anonymous):

Why does the base change to ten? I guess part of my confusion come from the fact that I don't quite understand what the base IS. Can you explain that?

OpenStudy (anonymous):

You can change in whatever base you like (10,2,5,6,7,8,25,32,.....) \[\log_{a} b=x=>a^x=b\]

OpenStudy (anonymous):

ex:\[\log_{2} 4=2\] because \[2^{2}=4\]

OpenStudy (anonymous):

Okay, i think I'm getting this. So, based on what you're saying, then that means that basically, logs are another form of exponential equations, right? But from what I read, logs are supposed to be reflections of exponential equations. Is that wrong?

OpenStudy (anonymous):

You're Right

OpenStudy (anonymous):

So then, how come \[\log _{a}b=x\] equals \[a^{x}=b\] Wouldn't they be the inverse of each other?

OpenStudy (anonymous):

you mean like\[a^{b}=x\]

OpenStudy (anonymous):

The equation you have in the fifth box up.

OpenStudy (anonymous):

The idea of logarithms is to reverse the operation of exponentiation, that is raising a number to a power.

OpenStudy (anonymous):

How is it reversed? Can you show me an example?

OpenStudy (anonymous):

\[Supposing:b^{y}=x=>\log_{b} x=y\]

OpenStudy (anonymous):

instead of y you have x in log

OpenStudy (anonymous):

Okay, I think I've got it now! It's making sense. Thank you for taking all that time!! I really appreciate it : )

OpenStudy (anonymous):

you're welcome